# Line of Sight Bias in Dark Matter Inferences from Galaxy Cluster Mergers

### v1

@@ Abstract
In collisions of galaxy clusters, the lack of displacement between dark matter and galaxies suggests that the dark matter scattering depth is small. This yields an upper limit on the dark matter cross section if the dark matter column density is known. We investigate a bias in such constraints: the measured column density (along the line of sight, using gravitational lensing) is lower than that experienced by a dark matter particle, as follows. Dark matter halos are triaxial and generally collide along their major axes, yielding a high scattering column density—but the merger is obvious only to observers whose line of sight is nearly perpendicular to that axis, yielding a low observed column density. We trace lines of sight through merging halos from the BigMDPL n-body simulation, both with and without mock observational effects. We find that a hypothetical skewer through the halo along the merger axis (more precisely, along the current separation vector of the two halos) has twice the column density of a typical line of sight. With weak lensing measurements, which involve some spatial averaging, this ratio is reduced to 1.25, suggesting that existing constraints on the scattering cross section are biased high by about 25%.
@@ 2 Methods
[…]
@@ 2.1 Geometric definitions
Figure 1: Definition of angles. $\psi$ is subtended by the halo major axis and the current separation vector to the other halo in the merger (not shown here). Lines of sight (LOS) fall along a cone defined by the separation vector and $\alpha$.
@@ 2.2 Simulation inputs and calculations
We used the publicly available Big Multidark Planck (BigMDPL, Klypin et al., 2016) simulation hosted on the CosmoSim [^1] https://www.cosmosim.org/metadata/bigmdpl/ website (Riebe et al., 2013). This dark matter simulation has a very large box size, $(2.5\ \textrm{Gpc}/h)^{3}$, which maximizes the number of merging halo pairs. The mass of each particle in the simulation is $2.359\times 10^{10}M_{\odot}$, so even the least massive clusters considered here ($M>6\times 10^{13}\,M_{\odot}$) have over 250 particles. The $\Lambda$CDM cosmological parameters were set to $h=0.6777$, $\Omega_{\Lambda}=0.692885$, and $\Omega_{m}=0.307115$.
We use the catalog of halo pairs assembled in Wittman et al. (2018a) and Wittman (2019). Briefly, this is derived from the BigMDPL Rockstar (Behroozi et al., 2013) halo catalog by finding halo pairs that have passed exactly one pericenter (within 300 kpc). This is in principle a different catalog at each snapshot, but we focus on one illustrative snapshot (number 74, $z=0.1058$). We discarded halo pairs with current separations $<0.25$ Mpc, as these were too overlapping for Rockstar to derive robust shape parameters. Observed mergers with current separations $<0.25$ Mpc are also rare because they are observationally difficult to disentangle.
@@
For weak lensing estimates, we use analytical formulae (Heyrovský & Karamazov, 2024) for the reduced shear field around a triaxial NFW Navarro et al. (1997) profile to generate mock weak-lensing data. We then fit the mock shear field, with a spherical NFW model, as is typically done in the observational literature. We then tabulate the ratio of inferred mass using a given LOS to that using the CSV.
@@ 3.1 Halo-merger alignment
For binary merging clusters specifically, we would expect good halo-halo alignment because both halos are falling toward each other along a filament with which they are each aligned. We are not aware of empirical measurements of this effect, which may be difficult given overlapping halos. However, halo-halo alignment in binary mergers can be inferred from two empirical results: the orientation of a brightest cluster galaxy (BCG) is a proxy for the orientation of its host cluster (West et al., 2017), and BCGs in merging clusters are known to align with the CSV (Wittman et al., 2019).
@@ 3.2 Skewer surface mass density
Illustrative halos. Figure 3, top panel, shows the results for an illustrative halo ($\psi=1.5^{\circ}$, $\frac{b}{a}=0.337$, $\frac{c}{a}=0.259$). Results are shown as violin plots for values of $\alpha$ in 5 increments. At $\alpha=0$, $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}=1$ by definition and only one LOS is sampled so no violin appears. As $\alpha$ increases, the the LOS pierces the halo closer to the intermediate or minor axis, so $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ falls below unity. Furthermore, at each $\alpha>0$ there are multiple LOS, yielding a range of $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ depending on their proximity to the halo minor or intermediate axis. Observed systems are typically idenitified as mergers and used for DM constraints only if $\alpha>45^{\circ}$, so this halo suggests that $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}\approx 0.3$. In other words, the surface mass density encountered by DM particles through the pericenter passage is about three times that measured along the LOS, so the resultant DM constraint would be biased high by a factor of three.
@@
Figure 3, bottom panel, shows a halo with matching axis ratios but with $\psi=18^{\circ}$ misalignment from the CSV. Therefore the maximum $\Sigma_{\rm los}$ is encountered by one LOS at $\alpha=18^{\circ}$: the LOS that happens to align with the major axis. Other LOS at $\alpha=18^{\circ}$ encounter a range of lower column densities. As one proceeds to higher $\alpha$ the LOS are sampling the intermediate and minor axes—but to a lesser degree than with a well-aligned halo, hence $\Sigma_{\rm los}$ does not drop as far.
All halos. A few halo pairs provided remarkably large or small bias factors, for example when one halo’s major axis points at the other’s minor axis in a “T” shape. We suspect that these extreme cases are artifacts of the halo finder having difficulty separating overlapping halos. To reduce clutter from these possibly unphysical cases, in Figure 4 we show the median and interquartile range (IQR) of the bias factor distribution after marginalizing over all halos in the relevant halo pairs. Again, considering that observed cases will have $\alpha>45^{\circ}$, the surface mass density encountered by DM particles through the pericenter passage is nearly twice that measured along the LOS, so the resultant DM constraint would be biased high by a factor of nearly two.
Figure 4: Median (circles) and interquartile range (bars) of the skewer bias factor marginalized over all halos, as a function of viewing angle $\alpha$. For typical observed cases $\alpha>45^{\circ}$ so the surface mass density along the LOS is about half that encountered by DM particles along the separation vector. At face value, this implies an overestimate of DM cross sections by a factor of two.
@@ 3.3 Mock weak lensing results
The skewer results presented so far represent a hypothetical measurement with unlimited angular resolution. Weak lensing (WL) involves a larger footprint on the sky so may be less sensitive to LOS. Our mock WL results confirm that, for a given halo and LOS pair, the mass ratio between the two LOS is less extreme for WL than for the skewer. At the largest $\alpha$, where a skewer along the LOS measures only half the surface density of a skewer along the CSV, we find that the WL ratio is 0.8. This more modest sensitivity to LOS is consistent with Euclid Collaboration et al. (2024) who performed mock WL fits along the principal axes of clusters in the Three Hundred simulation (Cui et al., 2018) and found that orientations along the (major, intermediate, minor) axes bias the inference of the true mass by factors of about (1.2, 0.85, 0.75). [^2] The Three Hundred halos include baryons so should be slightly rounder than our BigMDPL halos which are dark matter only. However, this is a small effect; see §4.
We show these results as a function of $\alpha$ in Figure 6. Again, the ratio is normalized to unity at $\alpha=0$, where there is a single LOS. At other $\alpha$ the median and interquartile range are shown, illustrating the variation across halos and LOS (e.g. LOS perpendicular to the major axis still sample a range from the minor to the intermediate axis). The full distribution is omitted for clarity. For large $\alpha$ the bias factor is just above 0.8.
Figure 6: Weak lensing bias factor as a function of viewing angle $\alpha$, marginalized over all halos. The circles indicate medians and the bars indicate the interquartile range. The trend is the same as for skewer measures of column density, but the bias factor is less extreme.
Figure 7 shows the same results split by mass quartile. For small to moderate $\alpha$, the WL results are consistent with the skewer results: low-mass halos are rounder so are less sensitive to LOS effects. However, at large $\alpha$ the low-mass halos actually exhibit more bias. The mechanism for this is unclear, as there are many steps in the WL fitting process. However, the pattern is consistent with Figure 10 of Euclid Collaboration et al. (2024), which found the inferred WL mass increasing more quickly with true mass when viewed along the intermediate and minor axes than when viewed along the major axis. In either case there is a great deal of scatter from halo to halo, so this trend is not highly significant.

### v2

@@ Abstract
In collisions of galaxy clusters, the lack of displacement between dark matter and galaxies suggests that the dark matter scattering depth is small. This yields an upper limit on the dark matter cross section if the dark matter column density is known. We investigate a bias in such constraints: the measured column density (along the line of sight, using gravitational lensing) is lower than that experienced by a dark matter particle, as follows. Dark matter halos are triaxial and generally collide along their major axes, yielding a high scattering column density—but the merger is obvious only to observers whose line of sight is nearly perpendicular to that axis, yielding a low observed column density. We trace lines of sight through merging halos from the BigMDPL n-body simulation, both with and without mock observational effects. We find that a hypothetical skewer through the center of the halo along the merger axis (more precisely, along the current separation vector of the two halos) has nearly twice the column density of a typical line of sight. With weak lensing measurements, which involve some spatial averaging, this ratio is reduced to 1.25, suggesting that existing constraints on the scattering cross section are biased high by about 25%.
@@ 2 Methods
We approximate each halo as a triaxial ellipsoid, meaning that its density $\rho$ is a function only of the scaled radial coordinate $r\equiv\sqrt{(\frac{x}{a})^{2}+(\frac{y}{b})^{2}+(\frac{z}{c})^{2}}$ where $a$, $b$, and $c$ denote the semiaxis lengths ($a>b>c$). We assume for notational simplicity here a Cartesian coordinate system aligned with the halo principal axes; in practice, this can be achieved by a rotation from the coordinate system of the simulation in which the halo is embedded.
2.1 Central Skewers
The surface mass density $\Sigma$ along any skewer piercing the center of the ellipsoid can be expressed as the skewer distance across the ellipsoid times an integral $\int_{0}^{\infty}\rho(r)dr$ that depends only on the density profile. Therefore the ratio $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ cancels the latter factor and involves only the ratio of distances across the ellipsoid. In this subsection we argue that this simple ratio is a useful thinking tool because it provides a bound on the size of the effect when integrated over larger surface areas, independent of the density profile or observational technique. Later in the paper, we perform more realistic estimates with mock weak gravitational lensing fits, which inherently average over a patch of sky.
Figure 1: Mean surface mass density along the minor axis, $\bar{\Sigma}_{\rm minor}$, as a fraction of that along the major axis, $\bar{\Sigma}_{\rm major}$, as a function of projected radius of the area enclosed, for a halo with minor axis ratio $\frac{c}{a}=0.4$. The central skewer prediction of 0.4 provides a lower bound independent of density profile.
Consider a LOS parallel to the major ($x$) axis, but displaced in $z$: Taylor expanding the density about $z=0$ produces a term proportional to $\frac{\rho^{\prime}(r)}{c^{2}}$. In contrast, expanding a minor-axis LOS about $x=0$ produces a term proportional to $\frac{\rho^{\prime}(r)}{a^{2}}$. Therefore, the density drops more quickly for LOS displaced from the major axis; including these non-central effects over a finite patch of sky will decrease $\bar{\Sigma}_{\rm major}$ faster than it decreases $\bar{\Sigma}_{\rm minor}$, where the bars indicate area averaging. Hence, skewers through the center provide a useful lower bound on the ratio of mean surface mass densities averaged over finite areas. We confirm this by using the formalism of Heyrovský & Karamazov (2024) to find $\frac{\bar{\Sigma}_{\rm minor}}{\bar{\Sigma}_{\rm major}}$ averaged over circles centered on the principal axes, for a variety of density profiles and circle sizes. Figure 1 shows the result for a halo with $\frac{b}{a}=0.6$ and $\frac{c}{a}=0.4$, for three profiles: NFW (Navarro et al., 1997), Einasto (Einasto, 1965), and Gaussian. The scale radius of the Einasto profile was chosen to match the NFW. The Gaussian profile is not physically motivated, but illustrates the behavior of an entirely different profile. With $\frac{c}{a}=0.4$, the central skewer argument predicts $\frac{\Sigma_{\rm minor}}{\Sigma_{\rm major}}=0.4$. Near the center, the Gaussian density profile is close to flat so the ratio remains close to 0.4 even when averaged over a substantial area. The other profiles are not flat near the center so $\frac{\bar{\Sigma}_{\rm minor}}{\bar{\Sigma}_{\rm major}}$ increases more rapidly with averaging radius. In all cases the central skewer ratio of 0.4 provides a lower bound, and overestimates the departure from unity by less than a factor of two.
Having established the utility of the central skewers, we now proceed to geometric definitions that will be used for both the skewer estimates and the mock weak lensing. At the end of this section we transition to describing the weak lensing simulations.
@@ 2.2 Geometric definitions
Figure 2: Geometric view. $\psi$ separates the halo major axis and the current separation vector to the other halo in the merger (not shown here). Lines of sight (LOS) fall along a cone defined by the separation vector and $\alpha$. At fixed $\psi$ and $\alpha$, LOS can vary substantially in column density.
@@ 2.3 Simulation inputs and calculations
[…]
We use the catalog of halo pairs assembled in Wittman et al. (2018a) and Wittman (2019). Briefly, this is derived from the BigMDPL Rockstar (Behroozi et al., 2013) halo catalog by finding halo pairs that have passed exactly one pericenter (within 300 kpc). This is in principle a different catalog at each snapshot, but we focus on one illustrative snapshot (number 74, redshift $0.1058$). We discarded halo pairs with current separations <0.25<0.25 Mpc, as these were too overlapping for Rockstar to derive robust shape parameters. Observed mergers with current separations <0.25<0.25 Mpc are also rare because they are observationally difficult to disentangle.
@@
The principal axes in the BigMDPL database are determined using the method of Allgood et al. (2006), which starts with a spherical region of radius of $R_{\rm vir}$ and calculates a reduced inertia tensor considering the particles within this region. Then, they update the region to an ellipsoidal region using the obtained axis ratios, keeping the semimajor axis fixed at $R_{\rm vir}$, and redo the calculation, iterating this process until it converges. This is a widely adopted robust way to determine axis ratios, and Bett (2012) showed that the median axis ratios produced this way are similar (within 0.05) to those yielded by a simple inertia tensor. We found that the median $\frac{c}{a}$ for BigMDPL halos above $0.6\times 10^{14}h^{-1}M_{\odot}$ is comparable to but slightly lower than those found in other studies: 0.46 vs. 0.45–0.50 for a similar mass range in Tenneti et al. (2015) and 0.55 for the ‘reduced iterated’ method in Bett (2012). The greater elongation of BigMDPL halos implies that we would find slightly smaller LOS bias effects in these other simulations. Still, it is not clear that BigMDPL halos are unduly elongated for our purposes given that halos are more elongated at higher density thresholds (Jing & Suto, 2002) where DM scattering effects will be most important. Our median $\frac{c}{a}$ of 0.46 agrees well with the median $\frac{c}{a}$ found in Jing & Suto (2002) for the high density threshold corresponding to the physical radius of 150 kpc we adopt below for averaging the surface mass density, even though the BigMDPL shapes were not explicitly evaluated at these high densities.
We checked the reliability of these axis ratios in a few ways. First, the BigMDPL database includes a second set of axis ratios, axisratio_[23]_1_500c, which are measured at a smaller radius than $0.3R_{\rm vir}$. We found that the statistics were quite similar, with median axis ratios differing only by $\approx 0.01$. Second, we found that in BigMDPL binary mergers, the axis ratios for the lower-mass partners match those of isolated halos (median $\frac{c}{a}=0.46$), but the more massive (’main’) partner is more elongated ($\frac{c}{a}=0.35$). This difference cannot be explained by particle misattribution near pericenter, which would presumably affect the lower-mass halo more. At least part of this difference may be the known effect that higher-mass halos are more elongated.
2.4 Weak Lensing
Inferences made via weak lensing average over a patch of sky and thus, unlike the central skewers, depend on the density profile. However, Figure 1 shows little difference between profiles commonly used to fit clusters. We therefore adopt the more commonly used NFW profile. We use the cataloged BigMDPL halo parameters m200c, rs, axisratio_2_1, and axisratio_3_1 to create a triaxial NFW model and fit it as follows. The workflow is summarized in Figure 3. Following Corless & King (2008), we use the spherical scale radius rs and the axis ratios to define a triaxial scale radius $R_{s}$, then use their Equation 7 to recover the triaxial $R_{200}$ given m200c. The ratio $\frac{R_{200}}{R_{s}}$ then forms a triaxial concentration which yields a scale density $\rho_{s}$. We then use the formalism of Heyrovský & Karamazov (2024) to obtain the true $\bar{\Sigma}_{\rm CSV}$ of the triaxial halo, which we store for reference. For this calculation, we use a fixed physical radius of 150 kpc as was used in the seminal dark matter constraint of Markevitch et al. (2004), although we vary this radius later to probe the robustness of the results.
Figure 3: Summary of the workflow for establishing a triaxial model halo and extracting inferences from mock observations.
We then choose a LOS and use the expressions in Heyrovský & Karamazov (2024) to generate a reduced shear field for that LOS. We discard any “observations” of shear in a 30′′ radius area around the center of the cluster, where cluster galaxies would prevent the detection of many background galaxies, and where the weak lensing approximation breaks down for massive clusters. To set the size of the simulated field, we tested a series of sizes with spherical halos and found that at 20 and larger diameter, the fitted mass converges to the true mass. We opted to use a 30 diameter field for the main run over all halos and all LOS.
To simulate the process of estimating $\bar{\Sigma}$ as it would typically be done in an observational paper, we then fit the mock shear field with a spherical NFW model following the Child et al. (2018) mass-concentration ($M$-$C$) relation, as weak lensing data are typically not informative enough to infer concentration or shape parameters. We fit with two free parameters, $M$ and $C$, which are linked by the $M$-$C$ relation but allowed to vary according to the scatter found by Child et al. (2018). In order to focus on the specific bias we are highlighting, we did not explore the effect of adding noise to the shear field. In that context, the central sky coordinates of the NFW halo were always well fit when allowed to float, so for efficiency in the main run we fixed those coordinates, leaving $M$ and $C$ as the only two free parameters.
For each halo and LOS, we use the best-fit spherical model to infer $\bar{\Sigma}_{\rm los}$ over a 150 kpc radius patch and normalize it by the known $\bar{\Sigma}_{\rm CSV}$ for that halo. In §3 we explore how $\frac{\bar{\Sigma}_{\rm los}}{\bar{\Sigma}_{\rm CSV}}$ varies as a function of viewing angle and halo properties.
@@ 3.1 Halo-merger alignment
For binary merging clusters specifically, we would expect good halo-halo alignment because both halos are falling toward each other along a filament with which they are each aligned. We are not aware of empirical measurements of this effect, which may be difficult given overlapping halos. However, halo-halo alignment in binary mergers can be inferred from two empirical results: the orientation of a brightest cluster galaxy (BCG) is a proxy for the orientation of its host cluster (West et al., 2017), and BCGs in merging clusters are known to align with the CSV (Wittman et al., 2019). Although we do not attempt a direct comparison because those works involve 2-D measurements of galaxies rather than 3-D measurements of clusters, the alignment seen in Figure 4 is about as strong as seen in those works.
@@ 3.2 Skewer surface mass density
Illustrative halos. Figure 5, top panel, illustrates how $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ varies with $\alpha$ for a very well aligned and very elongated halo ($\psi=1.5^{\circ}$, $\frac{b}{a}=0.337$, $\frac{c}{a}=0.259$). Results are shown as violin plots for values of $\alpha$ in 5 increments. At $\alpha=0$, $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}=1$ by definition and only one LOS is sampled so no violin appears. As $\alpha$ increases, the LOS pierces the halo closer to the intermediate or minor axis, so $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ falls below unity. Furthermore, at each $\alpha>0$ there are multiple LOS, yielding a range of $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ depending on their proximity to the halo minor or intermediate axis. Observed systems are typically identified as mergers and used for DM constraints only if $\alpha>45^{\circ}$, so this halo suggests that $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}\approx 0.3$. In other words, the surface mass density encountered by DM particles through the pericenter passage in this rather extreme case is about three times that measured along the LOS, so the resultant DM constraint would be biased high by a factor of three.
@@
Figure 5, bottom panel, shows an equally elongated halo with a more typical $\psi=18^{\circ}$ misalignment from the CSV. Therefore the maximum $\Sigma_{\rm los}$ is encountered by one LOS at $\alpha=18^{\circ}$: the LOS that happens to align with the major axis. Other LOS at $\alpha=18^{\circ}$ encounter a range of lower column densities. As one proceeds to higher $\alpha$ the LOS are sampling the intermediate and minor axes—but to a lesser degree than with a well-aligned halo, hence $\Sigma_{\rm los}$ does not drop as far. As a reminder, these are highly elongated halos chosen to illustrate the dependence of $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ on halo alignment $\psi$ and viewing angle $\alpha$. We now proceed to marginalize over all merging halos.
All halos. A few halo pairs provided remarkably large or small bias factors, for example when one halo’s major axis points at the other’s minor axis in a “T” shape. We suspect that these extreme cases are artifacts of the halo finder having difficulty separating overlapping halos. To reduce clutter from these possibly unphysical cases, in Figure 6 we show the median and interquartile range (IQR) of the bias factor distribution after marginalizing over all halos in the relevant halo pairs. Again, considering that observed cases will have $\alpha>45^{\circ}$, the surface mass density encountered by DM particles through the pericenter passage is nearly twice that measured along the LOS. Given that the central skewer estimates are a lower limit on $\frac{\bar{\Sigma}_{\rm los}}{\bar{\Sigma}_{\rm CSV}}$, we expect that resulting DM constraints would be biased high, but by less than a factor of two.
Figure 6: Median (circles) and interquartile range (bars) of the skewer bias factor marginalized over all halos, as a function of viewing angle $\alpha$. For typical observed cases $\alpha>45^{\circ}$ so the central surface mass density along the LOS is about half that encountered by DM particles along the separation vector. The gray curves illustrate the median behavior of each individual halo (medianized over various LOS at fixed $\alpha$).
@@ 3.3 Mock weak lensing results
We show the WL results as a function of $\alpha$ in Figure 8. Each halo is shown as a gray curve, with the medians and interquartile ranges overlaid. As expected, the results are not as extreme as the central skewer calculation: at large $\alpha$ the bias is about 0.8. This LOS bias happens to be of similar magnitude to that found by Euclid Collaboration et al. (2024) for total mass. They performed mock WL fits along the principal axes of clusters in the Three Hundred simulation (Cui et al., 2018) and found that orientations along the (major, intermediate, minor) axes bias the inference of the true mass by factors of about (1.2, 0.85, 0.75). [^2] The Three Hundred halos include baryons so should be slightly rounder than our BigMDPL halos which are dark matter only. However, this is a small effect; see §4. However, this is a numerical coincidence because our calculation is for surface mass density within 150 kpc projected radius rather than mass, and we show below that the result is sensitive to that particular choice.
Figure 8: Weak lensing bias factor as a function of viewing angle $\alpha$, marginalized over all halos. The circles indicate medians and the bars indicate the interquartile range. The gray curves illustrate the median behavior of each individual halo (medianized over various LOS at fixed $\alpha$). The trend is similar to that seen for the skewer estimate, but the bias factor is less extreme.
Note one distinction between the construction of the skewer and WL plots: for the skewers, the ratio $\frac{\Sigma_{\rm los}}{\Sigma_{\rm CSV}}$ was defined to be unity at $\alpha=0$. For the WL plots, observational effects and modeling choices could cause the inferred $\bar{\Sigma}_{\rm los}$ to differ from the true $\bar{\Sigma}_{\rm CSV}$ even when the LOS is along the CSV. (In fact, Figure 8 shows that the median ratio is slightly above unity at $\alpha=0$.) This choice in plot construction also forces the various halo curves to meet at $\alpha=0$ in the skewer plot (Figure 6) but not the WL plot (Figure 8). This causes an apparent variation in halo-to-halo scatter as a function of $\alpha$ in Figure 6.
Figure 9 shows the same results split by mass quartile. As expected, we find that the higher-mass halos are most sensitive to LOS effects, because they have more extreme axis ratios.
@@
Figure 10: Our fiducial weak lensing bias factor was defined in terms of surface mass density projected within 150 kpc of the halo center. Here we show how the bias factor depends on this choice.
To this point, the weak lensing bias factor has been defined in terms of surface mass density projected within 150 kpc of the halo center. Figure 10 shows the median bias (marginalized over all halos) as a function of $\alpha$ for two other choices of radius, 50 and 300 kpc. There is substantial variation, but mostly at small $\alpha$ which will rarely apply to observations in practice. At large $\alpha$, the variation is about $\pm 0.05$. This is comparable to the variation that might be seen with the use of different halo catalogs as discussed in §2.3.